OpenStudy (anonymous):

54 × 46 = (54 × 6) + ( × 40)

OpenStudy (kamibug):

\(\huge\cal\color{royalblue}{Welcome~to~Openstudy!~:)}\) Are you trying to solve for what number goes in the blank space?

OpenStudy (anonymous):

Idk

OpenStudy (kamibug):

Let's assume so. :P We can replace the blank space with a variable. Let's use the letter b. 54*46 = 54*6 + 40b First simplify the left side. What's 54*46?

OpenStudy (anonymous):

324

OpenStudy (kamibug):

That's 54*6, so now we have... 54*46 = 324 + 40b How about 54*46? :)

OpenStudy (anonymous):

2484

OpenStudy (nerdgirl):

will one of you guys help me with my problem when your done?

OpenStudy (kamibug):

Right! 2484 = 324 + 40b Now we'll combine like terms by subtracting 324 from both sides. 2484 - 324 = 324-324 + 40b 2484 - 324 = 0 + 40b 2484 - 324 = 40b What's 2484 - 324? :)

OpenStudy (anonymous):

2160

OpenStudy (anonymous):

so what go in the blank?

OpenStudy (anonymous):

idk

OpenStudy (kamibug):

Perfect. So the last step to solve for b is to divide both sides by 40. :] 2160/40 = b

OpenStudy (anonymous):

idk what that is

OpenStudy (anonymous):

idk the answer

OpenStudy (kamibug):

What's 2160 divided by 40? The quotient will be your answer.

OpenStudy (anonymous):

54

OpenStudy (kamibug):

You got it! ^_^

OpenStudy (anonymous):

Thanks!!! for ur help

OpenStudy (anonymous):

Can u help me with this to?16 + (10 – 8) ÷ 2

OpenStudy (anonymous):

wait sorry

OpenStudy (anonymous):

16+(10-8)/2

OpenStudy (kamibug):

Do what's in parenthesis first. What's 10-8?